# Sotopia-RL: Reward Design for Social Intelligence

### v1

@@ 4 Methodology
In this section, we introduce the details of our proposed Sotopia-RL framework. It includes two components: (1) social reward design with LLMs (2) social agent training with RL. The first part aims to provide high-quality offline reward labels for the training of utterance-level reward models, and the second part aims to use RL to optimize social agent policy.
@@ 4.1 Social Reward Design with LLMs
As discussed in Section 1, providing high‑quality utterance-level rewards for social interactions is challenging due to the features of social interactions. We simplify this reward design problem through two steps. First, rather than scoring utterances based on dialogue history, we attribute the episode‑level outcome to individual utterances based on the full episode, reducing variance in annotation. Second, we introduce a multi‑dimensional rubric that breaks down an utterance’s contribution into clearer aspects of goal achievement. Such decomposition converts a complex, noisy task into smaller, more structured judgments that LLMs can perform reliably, yielding fine‑grained and trustworthy supervision. Formally, given a social episode $\tau_{i}$, our goal is to assign each utterance $a_{t}^{i}$ an offline reward $r_{t}^{i}$ that reflects its contribution to the progression of the interaction.
Reward attribution: from episode‑level to utterance‑level. Episode‑level rewards offer only coarse supervision, since even conversations with low‑quality utterances can still lead to successful outcomes. Therefore, episode-level rewards $G_{i}$ are not an accurate estimate of utterance-level rewards for each $a_{t}^{i}$. To provide more precise feedback, we perform utterance‑level credit assignment. Advanced LLMs (e.g., GPT‑4o) evaluate each utterance within the full episode context, yielding attribution scores $\mathcal{A}_{i}(a_{t}^{i},\tau_{i})$ between 0 and 1 for each utterance $a_{t}^{i}$. Such offline attribution enables reliable credit assignment by leveraging global context. We then refine these scores using the episode‑level outcome $G_{i}$, so that strong utterances in successful episodes are emphasized, while good contributions in failed episodes remain acknowledged but proportionally down-weighted:
$r_{t}^{i}=G_{i}\cdot\mathcal{A}_{i}(a_{t}^{i},\tau_{i}).$ (2)
@@
Figure 4: Overview of social agent training. Our pipeline has three stages: (1) Data preparation: Generate GPT self‑play rollouts and annotate rewards with offline inference, where an LLM evaluates and attributes rewards after the conversation ends. (2) SFT training: Train the policy and an utterance‑level reward model via fine‑tuning. (3) RL training: Continue RL self‑play with rewards from online inference, where the RM sees only the dialogue history up to the current turn for prediction.
Reward combination: from single to multi-dimension. While utterance‑level reward attribution improves granularity, attributing to a single reward $G_{i}$ alone cannot capture the quality of an utterance. High‑quality utterances not only advance the agent’s goal but also foster engagement and sustain conversational flow. To capture such broader contributions, we extend rewards beyond the primary goal completion score $G_{i}$ to additional dimensions from Sotopia-eval. Specifically, relationship maintenance (Rel) assesses whether an interaction preserves or strengthens the relationship between agents, and knowledge seeking (Kno) assesses whether two agents gain new knowledge during interactions (full definitions in Appendix §H). These two dimensions are included for attribution because they promote collaborative and information‑rich interactions and indirectly contribute to goal achievement, whereas other metrics in Sotopia-eval are often too task‑specific to generalize. Therefore, for the utterance $t$ of agent $i$, we collect attributed dimension‑wise scores $r_{t,d}^{i}$ based on Eq. 2, normalize them within each dimension, and aggregate them into a final reward:
$r_{t}^{i}=\frac{1}{N}\sum_{d=1}^{N}\gamma_{d}\cdot\frac{r_{t,d}^{i}-\min(r_{\cdot,d}^{i})}{\max(r_{\cdot,d}^{i})-\min(r_{\cdot,d}^{i})}.$ (3)
Here $N$ is the number of reward dimensions, $\gamma_{d}$ is the weight for dimension $d$, and $\min(r_{\cdot,d}),\max(r_{\cdot,d})$ denote the range of scores for that dimension across the dataset.
Overall: offline and flexible reward design. As shown in Figure 4, we begin with a single episode‑level reward and expand it along two axes to obtain denser supervision. This process yields utterance‑level multi-dimensional rewards that capture the quality of each contribution. The design is offline because it leverages the full dialogue context after the episode concludes, allowing for better evaluation without the uncertainty of real‑time inference. It is flexible because the rubric can be extended with additional dimensions and easily adapted to different social contexts—for example, prioritizing relationship building (Rel) in therapeutic dialogue or focusing on knowledge seeking (Kno) in educational tutoring.
@@ 4.2 Social Agent Training with RL
As shown in Figure 4, Stage 1 is about social reward design, and in Stages 2 and 3, our key contribution is bridging the POMDP nature of social interactions with the scalability of MDP‑based RL through RM distillation. By distilling the knowledge in offline rewards into an utterance-level RM that provides online utterance‑level feedback, we enable efficient training while teaching the policy to implicitly infer about the latent social context.
Training utterance-level reward models. Directly applying POMDPs to LLM training is impractical because belief tracking over high‑dimensional latent states in social interactions is computationally intractable, making MDP‑based RL with RM distillation the scalable alternative. As shown on the right of Figure 4, we therefore compress offline reward signals into an RM that provides utterance‑level feedback conditioned on dialogue history. While the MDP approximation alone risks overlooking latent social factors that are only partially observable in the local dialogue history $s_{t}^{i}$, our RM $R_{\psi}(s_{t}^{i},a_{t}^{i})$ is trained to recover these signals by distilling the offline rewards into a function that infers hidden context from local observations. We train $R_{\psi}$ to regress towards the offline rewards $r_{t}^{i}$ through
$\mathcal{L}_{\text{MSE}}=\mathbb{E}_{(s_{t}^{i},a_{t}^{i})}\left[\left(R_{\psi}(s_{t}^{i},a_{t}^{i})-r_{t}^{i}\right)^{2}\right].$ (4)
This allows the resulting RM $R_{\psi}$ to provide accurate, fine‑grained feedback for RL optimization within the tractable MDP framework, while still capturing the latent dimensions of social quality that a full POMDP would model.
Training policy models with single-turn online RL. As shown on the left of Figure 4, we train the social agent policy $\pi_{\theta}$ with a reward model $R_{\psi}$ that provides utterance-level feedback. Training proceeds in two stages. First, we warm up the policy with behavior cloning (BC) on GPT self-play rollouts to ensure coherent generation. We then further train the policy with GRPO (Shao et al., 2024b), adopting a single-turn formulation for efficiency. Each self-play rollout is decomposed into multiple $(s_{t}^{i},a_{t}^{i})$ pairs, and at each step $t$ the policy generates $a_{t}^{i}\sim\pi_{\theta}(\cdot\mid s_{t}^{i})$, receives a reward $R_{\psi}(s_{t}^{i},a_{t}^{i})$, and updates to maximize expected rewards. We do not add explicit reasoning traces during inference, focusing on optimizing utterances efficiently. Although simplified to single-turn updates, the distillation from offline rewards allows the agent to achieve improved performance in multi-turn interactions.
5 Experimental Settings
Model settings. We select Qwen2.5-7b-Instruct  [^1] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct as our base LLM for the training of both the policy model and reward model. We select GPT-4o for LLM-as-the-judge in Sotopia-eval.
@@ 5 Experimental Settings
Goal Overall Goal Overall
@@
Table 1: Our method outperforms state-of-the-art models when choosing GPT-4o as partner ($p<0.05$, paired t-test on the Goal dimension). Qwen2.5-7B refers to Qwen-2.5-7B-Instruct. Training method baselines from PPDPP to DSI have details available in Section §5. Full experimental results are available in Appendix §E.1.
@@
Attribution Dimension Sotopia-hard
Rel Kno Goal Overall
Qwen2.5-7B Uniform Goal 1.84 4.14 5.61 2.95
Singular Goal 2.74 4.93 6.64 3.41
Scaled Goal 1.82 3.83 6.74 3.15
Direct Rel 3.61 4.92 7.24 3.60
Direct Kno 2.56 6.06 6.93 3.61
Direct Goal 2.49 4.94 7.21 3.49
+Behavior Cloning (BC) 2.49 3.37 6.76 3.16
+Sotopia-$\pi$ 2.41 3.66 6.84 3.20
+Sotopia-RL 3.41 5.53 7.81 3.80
@@
Table 2: Our social reward designs outperform reward design baselines with the BC model as partner ($p<0.05$, paired t-test on the Goal dimension). All results use Qwen2.5-7B-Instruct as the base model. Goal-RL refers to direct+Goal; Sotopia-RL combines Rel, Kno, and Goal via direct attribution by averaging. BC and Sotopia-$\pi$ are baselines without reward design. Full experimental results are provided in Appendix §E.1.
@@
Evaluation settings. We evaluate our method on two configurations of the Sotopia benchmark: (1) Sotopia-hard, and (2) Sotopia-all. Sotopia-hard is a subset of Sotopia-all, consisting of 14 challenging social scenarios identified as difficult among all scenarios, and we use 10 distinct agent pairings per scenario. For Sotopia-all, we evaluate on the full coverage of 90 social scenarios, using 2 agent combos per scenario to ensure diversity while maintaining scalability. More statistical information is in Appendix §A.
@@
Training method baselines. To compare the effectiveness of our training methods, we include (1) behavior cloning (BC) that utilizes social interaction trajectories between GPT-4o, which is the same as Sotopia-$\pi$; (2) Sotopia-$\pi$ (Wang et al., 2024b) that utilizes behavior cloning and self-reinforcement; (3) other most recent baselines: PPDPP (Deng et al., 2024), EPO (Liu et al., 2025), DAT (Li et al., 2024), and DSI (Zhang et al., 2025). Sotopia-RL denotes our proposed approach, which combines direct utterance-level attribution with a multi-dimensional reward combination design (Rel +Kno +Goal), trained using single-turn online RL (GRPO) without explicit reasoning but just utterance generation. Training details are available in Appendix §B.
@@
Reward attribution baselines. To assess the effectiveness of our reward attribution strategy, we compare it against four baselines, each defining how utterance-level rewards $r_{t}^{i}$ are derived from the episode-level score $G_{i}$: (1) Uniform — every utterance receives the same reward, $r_{t}^{i}=G_{i}$ for all $t$; (2) Singular — only one selected utterance $a_{k}^{i}$ is assigned the full reward, $r_{t}^{i}=G_{i}$ if $t=k$ and $r_{t}^{i}=0$ otherwise; (3) Scaled — the episode-level reward is distributed proportionally, $r_{t}^{i}=\alpha_{t}G_{i}$ with $\sum_{t}\alpha_{t}=1$ and $\alpha_{t}\geq 0$; and (4) Direct — each utterance is independently attributed with a normalized weight, $r_{t}^{i}=\alpha_{t}G_{i}$, where each $\alpha_{t}\in[0,1]$ reflects its contextual attribution score, ensuring no utterance exceeds the episode-level score and same with Eq. 2. Direct attribution is the method used in Sotopia-RL. Details for each attribution method are in Appendix §G.
Reward combination baselines. We train our RM with labels combining three Sotopia-eval dimensions: relationship maintenance (Rel), knowledge seeking (Kno), and goal completion (Goal). Formally, the single-dimension baselines use $r_{t}^{i}=r_{t,\textsc{Rel}}^{i}$, $r_{t}^{i}=r_{t,\textsc{Kno}}^{i}$, or $r_{t}^{i}=r_{t,\textsc{Goal}}^{i}$, corresponding to Rel-only, Kno-only, and Goal-only for Eq. 3, respectively. Sotopia-RL averages the three signals, $r_{t}^{i}=\tfrac{1}{3}\big{(}r_{t,\textsc{Rel}}^{i}+r_{t,\textsc{Kno}}^{i}+r_{t,\textsc{Goal}}^{i}\big{)}$ for Eq. 3, allowing us to isolate the contribution of each component while validating the benefit of multi-dimensional reward modeling. We use a simple average to give equal importance to each reward dimension. Conducting a more fine‑grained weighting would require hyper‑parameter search, which is computationally expensive in the Sotopia environment. See Appendix §H for details.
@@ 6 Experimental Results
Sotopia-RL helps build the state-of-the-art social agents on the Sotopia benchmark. In Table 1, Qwen-2.5-7B-Instruct trained with Sotopia-RL reaches the highest goal completion score, achieving the 7.17 score in the Sotopia-hard. It indicates that our utterance-level RM provides better guidance during the training of RL. It also indicates that for multi-turn social interactions, improving the quality of single-turn interactions with suitable single-turn rewards can effectively optimize multi-turn performance. Notably, AMPO (Wang et al., 2025) reaches 7.50 on Sotopia-hard. But it includes an explicit reasoning process and requires more than 640 inference tokens per utterance on average. Therefore, it is unfair to compare AMPO with ours since we only utilize GRPO to generate utterances without extra tokens for reasoning. Full results are available in Appendix §E.
@@
Figure 5: Evaluation results with different LLM-based evaluators. The consistent improvement on evaluators indicates no reward hacking. Full results in Appendix §E.2.
Figure 6: Evaluation results with different partner models. The consistent improvement with multiple partners indicates no reward hacking. Full results in Appendix §E.2.
@@
Reward combination contributes to the performance improvement. Based on Table 2, training with multiple reward dimensions—goal completion (Goal), relationship maintenance (Rel), and knowledge seeking (Kno)—significantly improves performance compared to using a single reward dimension. The best result comes from combining all three dimensions, yielding a 7.9This improvement arises because multi‑objective RL encourages the policy to balance different aspects of interaction quality when generating utterances. Notably, as shown in Table 2, optimizing each dimension independently also improves performance on the others, suggesting that the objectives are correlated. Thus, combining them makes reward model training more robust.

### v3

@@ 4 Methodology
In Figure 4, we introduce the proposed Sotopia-RL framework, which consists of two stages: (1) offline social reward design with LLMs and (2) online social agent training with RL. In the offline stage, high-quality reward labels are generated using the entire dialogue episode, where each utterance is annotated by LLMs with access to both its preceding and subsequent context. These offline labels are then distilled into an online reward model (RM) that relies only on the dialogue history available up to the current utterance. In the online stage, this RM is used to provide real-time reward signals during interaction, enabling policy optimization through online RL.
@@ 4.1 Social Reward Design with LLMs
Providing accurate utterance-level rewards for social interactions is challenging due to the uniqueness of social intelligence tasks mentioned in Section §1. We address this challenge in two steps, as shown in Figure 4. First, through reward attribution, we assign episode-level outcomes to individual utterances based on the full dialogue, rather than scoring them only from local context. This reduces variance in annotation and provides more stable supervision. Second, through reward aggregation, we introduce a multi-dimensional rubric that decomposes an utterance’s contribution into distinct aspects of goal achievement. This decomposition turns a complex and noisy evaluation problem into smaller, structured judgments that LLMs can perform more reliably. Formally, given a social episode $\tau$, our goal is to assign each utterance $a_{t}$ an offline reward $r_{t}$ that reflects its contribution to the progression of the interaction.
Reward attribution: from episode-level to utterance-level. Episode-level rewards provide only coarse supervision, since individual social utterances are only weakly correlated with final success. This makes RL training on episode-level signals unstable and inefficient. Consequently, episode-level rewards $G$ are not accurate estimates of the true contribution of each utterance $a_{t}$. To provide more fine-grained feedback, we perform utterance-level credit assignment. Advanced LLMs (e.g., GPT-4o) evaluate each utterance within the full episode context, producing attribution scores $\mathcal{A}(a_{t},\tau)\in[0,1]$. This offline attribution leverages global context to assign credit more reliably. We then refine these scores with the episode-level outcome $G$, ensuring that strong utterances in successful episodes are emphasized, while contributions in failed episodes are still recognized but proportionally down-weighted:
$r_{t}=G\cdot\mathcal{A}(a_{t},\tau).$ (3)
Figure 3: Overview of the Sotopia-RL pipeline. Stage 1—Offline data preparation: generate GPT self-play dialogues and assign utterance-level rewards via offline inference with full-episode context. Stage 2—RL training: (2.1) SFT initializes the policy and distills an utterance-level reward model (RM) from offline labels; (2.2) Online RL method continues self-play using rewards from the online RM, which conditions only on dialogue history up to the current turn.
@@
Reward aggregation: from single to multi-dimension. While utterance-level reward attribution improves granularity, relying solely on a single goal-completion score $G$ cannot fully capture utterance quality. High-quality utterances in a dialogue are not always tied directly to the goal itself—some utterances maintain engagement, sustain conversational flow, or strengthen social bonds, which are equally important for successful interactions. Simply optimizing toward goal completion at every turn risks overemphasizing short-term task progress while neglecting these broader aspects of social interaction. To address this limitation, we incorporate all seven evaluation dimensions from Sotopia as the rubric for reward design. For each utterance, the LLM evaluator provides attributed scores along these dimensions using the same attribution procedure (Eq. 3). We then normalize scores within each dimension and aggregate them into a final reward through a weighted average over $N$ dimensions:
$\tilde{r}_{t,d}=\frac{r_{t,d}-\min\nolimits_{k}r_{k,d}}{\max\nolimits_{k}r_{k,d}-\min\nolimits_{k}r_{k,d}},\qquad r_{t}=\frac{1}{N}\sum_{d=1}^{N}\gamma_{d}\cdot\tilde{r}_{t,d}.$ (4)
Here $\tilde{r}_{t,d}$ denotes the normalized reward for dimension $d$ at time $t$ and $\gamma_{d}$ is its corresponding weight. Empirically, we find that among all dimensions, relationship maintenance (Rel) and knowledge seeking (Kno) play a particularly crucial role for better goal achievement. It is potentially because these dimensions help build better conversational flow.
Overall: offline and rubric-based reward design. As shown in Figure 4, we start from a single episode-level reward $G$ and examine the limitations of using it directly as the reward signal. To address these issues, we expand the reward along two axes to provide denser supervision, ultimately yielding utterance-level, multi-dimensional rewards that capture the quality of each utterance. The design is offline, as it leverages the full dialogue context once the episode concludes, enabling more reliable evaluation without the uncertainty of real-time inference. It is also rubric-based, since additional dimensions can be easily incorporated and tailored to different social contexts—for example, prioritizing relationship building (Rel) in therapeutic dialogue or emphasizing knowledge seeking (Kno) in educational tutoring.
@@ 4.2 Social Agent Training with RL
After collecting utterance-level and multi-dimensional reward labels, the next step is to leverage these rewards for RL training. The key challenge here is to construct a reliable reward model (RM) that can estimate the quality of an utterance solely from the preceding dialogue context. Such a model not only provides consistent reward signals but also enables online RL training.
Training utterance-level multi-dimensional reward models. To bridge offline labels with online training, we distill the global information from full episodes into an utterance-level RM. For each dialogue turn, represented by the state–action pair $(s_{t},a_{t})$ consisting of the current social interaction context $s_{t}$ and the candidate utterance $a_{t}$, the RM is trained to anticipate the potential outcomes and assign an appropriate score. Supervised by the designed reward $r_{t}$, it outputs a scalar reflecting utterance quality. The model is optimized using mean squared error (MSE) loss:
$\mathcal{L}_{\text{MSE}}=\mathbb{E}_{(s_{t},a_{t})}\left[\left(R_{\theta}(s_{t},a_{t})-r_{t}\right)^{2}\right].$ (5)
This enables the trained RM $R_{\theta}$ to translate rich, multi-dimensional feedback into accurate turn-level signals, supporting stable and fine-grained RL optimization.
Training policy models with single-turn online RL. Given a trained online reward model, we can now optimize the social policy using standard RL algorithms. As shown on the left of Figure 4, the social agent policy $\pi_{\theta}$ is trained with the reward model $R_{\theta}$, which provides utterance-level feedback. Training proceeds in two stages. First, we warm up the policy with behavior cloning (BC) on GPT self-play rollouts to establish coherent generation. We then fine-tune the policy with GRPO (Shao et al., 2024b), adopting a single-turn formulation for efficiency. In this setup, each self-play rollout is decomposed into multiple $(s_{t},a_{t})$ pairs. At each step $t$, the policy generates $a_{t}\sim\pi_{\theta}(\cdot\mid s_{t})$, receives a reward $R_{\theta}(s_{t},a_{t})$, and updates its parameters to maximize expected rewards. We deliberately avoid adding explicit reasoning traces during inference, focusing instead on efficient utterance optimization. Although simplified to single-turn updates, the distilled multi-turn information from RM equips the policy to perform effectively in multi-turn social interactions.
5 Experimental Setup
Model settings. We select Qwen2.5-7b-Instruct [^1] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct as our base LLM for the training of both the policy model and reward model. We select GPT-4o [^2] https://platform.openai.com/docs/models/gpt-4o for LLM-as-the-judge in Sotopia.
@@ 5 Experimental Setup
Goal Avg Goal Avg
@@
Table 1: Our method outperforms state-of-the-art models when choosing GPT-4o as partner ($p<0.05$, paired t-test on the Goal dimension). Qwen2.5-7B refers to Qwen-2.5-7B-Instruct. Training method baselines from PPDPP to DSI have details available in Section §5. Full experimental results are available in Appendix §A.1.
@@
Reward Attribution Reward Dimension Sotopia-hard
Bel Rel Kno Goal Avg
Qwen2.5-7B Uniform Goal 8.81 1.84 4.14 5.61 2.95
Singular Goal 9.00 2.74 4.93 6.64 3.41
Scaled Goal 8.94 1.82 3.83 6.74 3.15
Direct Bel 8.98 2.66 4.56 6.93 3.37
Direct Rel 8.96 3.61 4.92 7.24 3.60
Direct Kno 8.99 2.56 6.06 6.93 3.61
Direct Goal 8.99 2.49 4.94 7.21 3.49
+Behavior Cloning (BC) 9.01 2.49 3.37 6.76 3.16
+Sotopia-$\pi$ 8.99 2.41 3.66 6.84 3.20
+Sotopia-RL 9.01 3.41 5.53 7.81 3.80
@@
Table 2: Our social reward designs outperform reward design baselines with the BC model as partner ($p<0.05$, paired t-test on the Goal dimension). All results use Qwen2.5-7B-Instruct as the base model. Goal-RL refers to Direct+Goal; Sotopia-RL combines Rel, Kno, and Goal via Direct attribution by averaging. BC and Sotopia-$\pi$ are baselines without reward design. Full experimental results are provided in Appendix §A.1.
@@
Evaluation settings. We evaluate our method on two configurations of the Sotopia benchmark: (1) Sotopia-hard, and (2) Sotopia-all. Sotopia-hard is a subset of Sotopia-all, consisting of 14 challenging social scenarios identified as difficult among all scenarios, and we use 10 distinct agent pairings per scenario. For Sotopia-all, we evaluate on the full coverage of 90 social scenarios, using 2 agent combos per scenario to ensure diversity while maintaining scalability. More statistics about the dataset are in Appendix §D. We report evaluation metrics on believability (Bel), relationship building (Rel), knowledge seeking (Kno), goal completion (Goal), and overall average score (Avg). More details about evaluation dimensions are in Appendix §I.
@@
Training method baselines. To compare the effectiveness of our training methods, we include (1) behavior cloning (BC) that utilizes social interaction trajectories between GPT-4o, which is the same as Sotopia-$\pi$; (2) Sotopia-$\pi$ (Wang et al., 2024b) that utilizes behavior cloning and self-reinforcement; (3) other most recent baselines: PPDPP (Deng et al., 2024), EPO (Liu et al., 2025), DAT (Li et al., 2024), and DSI (Zhang et al., 2025). Sotopia-RL denotes our proposed approach, which combines direct utterance-level attribution with a multi-dimensional reward aggregation design (Rel +Kno +Goal), trained using single-turn online RL (GRPO) without explicit reasoning but utterance generation. Training details are available in Appendix §E.
@@
Reward attribution baselines. To assess the effectiveness of our reward attribution strategy, we compare it against four baselines, each defining how utterance-level rewards $r_{t}$ are derived from the episode-level score $G$: (1) Uniform — every utterance receives the same reward, $r_{t}=G$ for all $t$; (2) Singular — only one selected utterance $a_{k}$ is assigned the full reward, $r_{t}=G$ if $t=k$ and $r_{t}=0$ otherwise; (3) Scaled — the episode-level reward is distributed proportionally, $r_{t}=\alpha_{t}G$ with $\sum_{t}\alpha_{t}=1$ and $\alpha_{t}\geq 0$; and (4) Direct — each utterance is independently attributed with a normalized weight, $r_{t}=\alpha_{t}G$, where each $\alpha_{t}\in[0,1]$ reflects its contextual attribution score, ensuring no utterance exceeds the episode-level score and same with Eq. 3. Direct attribution is the method used in Sotopia-RL. Details for each attribution method are in Appendix §H.
Reward aggregation baselines. To assess the effectiveness of our reward aggregation, we include four single-dimension baselines: Bel-RL ($r_{t}=r_{t,\textsc{Bel}}$), Goal-RL ($r_{t}=r_{t,\textsc{Goal}}$), Kno-RL ($r_{t}=r_{t,\textsc{Kno}}$), and Rel-RL ($r_{t}=r_{t,\textsc{Rel}}$). Sotopia-RL selectively average three signals as in Eq. 4:
$r_{t}=\tfrac{1}{3}\big(r_{t,\textsc{Rel}}+r_{t,\textsc{Kno}}+r_{t,\textsc{Goal}}\big),$
This setup allows us to isolate the contribution of each component while confirming the benefits of multi-dimensional reward modeling. We use a simple average to treat all dimensions equally, as our ablation study shows that the choice of weights has little impact on performance. Full experiments on dimension ablation and weight ablation are in Appendix §I.
@@ 6 Experimental Results
Sotopia-RL helps build state-of-the-art social agents on the Sotopia benchmark. In Table 1, Qwen-2.5-7B-Instruct trained with Sotopia-RL reaches the highest goal completion score, achieving the 7.17 score in the Sotopia-hard. It indicates that our utterance-level RM provides better guidance during the training of RL. It also indicates that for multi-turn social interactions, improving the quality of single-turn interactions with suitable single-turn rewards can effectively optimize multi-turn performance. Notably, AMPO (Wang et al., 2025) reaches 7.50 on Sotopia-hard. But it includes an explicit reasoning process and requires more than 640 inference tokens per utterance on average. Therefore, it is unfair to compare AMPO with ours since we only utilize GRPO to generate utterances without extra tokens for reasoning. Full results are available in Appendix §A.
@@
Figure 5: Evaluation results with different LLM-based evaluators. The consistent improvement on evaluators indicates no reward hacking. Full results in Appendix §A.2.
Figure 6: Evaluation results with different partner models. The consistent improvement with multiple partners indicates no reward hacking. Full results in Appendix §A.2.
@@
Reward aggregation contributes to the performance improvement. Based on Table 2, training with multiple reward dimensions—goal completion (Goal), relationship maintenance (Rel), and knowledge seeking (Kno)—significantly improves performance compared to using a single reward dimension. The best result comes from combining all three dimensions, yielding a 7.9% gain in goal completion (Goal). This improvement arises because multi‑objective RL encourages the policy to balance different aspects of interaction quality when generating utterances. Notably, as shown in Table 2, optimizing each dimension independently also improves performance on the others, suggesting that the objectives are correlated. Thus, combining them makes RM training more robust.